OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

-2(x+3)=-2(x+1)-4 -2x-6=-2x-1-4 -2x+2x=-5+6 No solution thus -2x+2x=0 :)